find the equation
a line perpendicular to y=3x-2 and passing through (2,3)

Answer by Alan3354(69443)

find the equation of a line perpendicular to y=3x-2 and passing through (2,3)
----------------------
1st find the slope of y=3x-2
Since it's in the slope-intercept form, y = mx+b, the slope (m) is 3.
Lines perpendicular have a slope of the negative inverse, or -1/3.
Use the slope-intercept again:
y-y1 = m*(x-x1) where (x1,y1) is (2,3)
y-3 = (-1/3)*(x-2)
3y-9 = -x+2
x + 3y = 11 (standard form)
y = (-1/3)x + 11/3 (slope-intercept form)